OpenStudy (hippoharry):

A set of data points has an equation of = -1.5 + 2.5x. What is the residual for the point (2, 5)? A. 3.5 B. 5 C. -1.5 D. 1.5

OpenStudy (hippoharry):

that doesnt make sense

Vocaloid (vocaloid):

x = 2 means you can replace "x" with "2" whenever x shows up

Vocaloid (vocaloid):

now find y

OpenStudy (hippoharry):

oh okay i understand

Vocaloid (vocaloid):

y = -1.5 + 2.5x.

Vocaloid (vocaloid):

what is y?

OpenStudy (hippoharry):

wait so i plugged in 2 for x and then got y=-1.5-5

OpenStudy (hippoharry):

then i got -6.5

OpenStudy (hippoharry):

im still confused

Vocaloid (vocaloid):

i made a typo in the first equation

Vocaloid (vocaloid):

y = -1.5 + 2.5x. = ?

OpenStudy (hippoharry):

what do i solve for @Vocaloid

Vocaloid (vocaloid):

y

Vocaloid (vocaloid):

what is y when x = 2?

OpenStudy (hippoharry):

i got 3.5

OpenStudy (hippoharry):

is that correct?

Vocaloid (vocaloid):

yes now, this value tells us the "predicted" value the point (2,5) from the original problem is the "actual" value (y = 5) so, to find the residual we do actual - predicted = 5 - 3.5 = 1.5

Vocaloid (vocaloid):

and that's it